# arXiv:2404.05610v1[cs.DC] 08 Apr 2024

### 08 Apr 2024 v1

@@ III-B Input and Output Parameters
Since one is primarily interested in the receive buffer in MPI calls, this parameter is always implicitly returned by KaMPIng. To retrieve other parameters from the wrapped MPI call one has to explicitly pass the corresponding out-parameter. This is a major improvement over previous MPI libraries, which simply mimic MPI’s C-Interface and return output data next to the receive buffer by pointer or reference. This is not in line with the C++core guidelines which strongly suggest to return output data by value [16, F.20]. Furthermore, it is often unclear which additional parameters are computed by the library, as the overloaded wrapped MPI function calls differ only in the number of function arguments. Combined with the large number of parameters of MPI calls, it is hard to see which argument refers to which parameter when looking at the code.
@@
In the following this is illustrated with a call to the wrapped MPI_Allgatherv function.
@@
A shortcoming of returning by value is the redundant memory allocation in case a previously allocated container could be reused instead. In KaMPIng, we offer two solutions for this scenario. For containers supporting C++move semantics, a previously user allocated container can be simply moved to the underlying call and is then subsequently returned with the result object by value.
@@
If there is no efficient way to support move semantics for a container type, it is also possible to pass the container via reference to the underlying call. The data computed by KaMPIng for the requested parameter will then be written directly to the specified memory-location.
@@ III-C Controlling Memory Allocation
Previous MPI wrappers have no unified way of controlling memory allocation. They either accept containers which are always resized to fit, or, if resizing is not desired, the user has to pass raw pointers directly. They also offer no control over allocation happening for default parameter computation.
@@
If KaMPIng has to create auxiliary data structures to compute missing parameters, the user may either pass preallocated containers to use or can provide the container’s type via template parameters. Recall that additional allocation is omitted entirely when parameters are provided by the user.
As all of this relies on template meta-programming, there is no additional overhead compared to a hand-rolled implementation. This flexibility allows to quickly implement distributed algorithms and then iteratively fine-tune memory allocations and library inferred values. This facilitates an algorithm engineering workflow which involves iterative refinement of implementations and analysis through experimentation.
@@ III-D Using Custom Types
HPC applications use a variety of data types which need to be communicated. Beyond basic data types corresponding directly to C++’s built-in types, MPI allows for complex derived data types using type constructors, such as MPI_Type_create_struct and MPI_Type_create_contiguous.
C’s lack of type-introspection forces users to always pass the type explicitly to a communication call, which is both tedious and error-prone, as type definitions need to be kept in sync with the actual data layout.
@@ III-D1 Static derived data types at compile time
namespace __mping {
// using __MPIng’s built-in struct serializer
@@
Existing MPI bindings also support non-built-in static types to some extent. For Boost.MPI, this is dependent on the definition of a serialization function. Similar to KaMPIng, types are managed using a global type pool, but each operation incurs a runtime type lookup. MPL and RWTH-MPI use the Construct-On-First-Use-Idiom to commit types before first use, and MPL also uses PFR to provide automatic type definitions by using reflection. Opposed to KaMPIng, types are not properly freed, which may result in resource leakage. RWTH-MPI allows custom static type definitions, but the user is responsible for committing and freeing types.
@@ III-D2 Support for dynamic types
In KaMPIng, dynamic types can currently only be constructed using MPI’s type constructors. MPL on the other hand provides a runtime type interface mirroring MPI’s type constructors using the builder pattern, which encapsulates the constructed types in so-called layouts. While MPL’s collective operations are tightly interleaved with the layout system, which results in verbose code, its type construction is a powerful feature which we plan to integrate as the default way for constructing dynamic types in KaMPIng. RWTH-MPI offers no support for dynamic types.
@@ III-D3 Communicating arbitrary data using serialization
Some applications require communicating non-contiguous data which is (partially) located on the heap, e.g., std::string or std::unordered_map. These have to be packed to a contiguous buffer before communication. KaMPIng facilitates this by providing serialization support, which is both highly tuneable and transparent to the user. We rely on the popular C++serialization library Cereal [19], which supports STL data types and allows providing serialization routines for custom types. While serialization is transparent to the user, it also has to be explicitly enabled as it usually incurs hidden costs for allocation memory for serialized data and perform (de-)serialization. Through Cereal’s flexible design, serialization in KaMPIng is also highly configurable; allowing users to specify custom serialization functions and archives, e.g., binary formats, JSON, or XML. See Figure 3 for an example on how to use serialization in KaMPIng.
@@
send_buf(as_deserializable<dict>())
@@
Besides KaMPIng, Boost.MPI is the only library offering serialization support, but is tightly coupled with other Boost libraries and serialization is performed implicitly; if a type is not marked as an MPI data type, serialization is used.
@@ III-D4 Towards sensible defaults for type construction
By default, KaMPIng maps trivially-copyable types to a type interpreted as a contiguous sequence of bytes. When defining a struct type where the members have alignment gaps, MPI does not include the gap in the communicated data. This requires non-contiguous memory accesses, which may be slower than copying whole memory blocks to the communication hardware (MPI standard [1, §5.1.6]). The standard suggest introducing dummy struct members to fill these gaps, but this requires the user to modify their non-MPI code. By using a type consisting of contiguous bytes when valid with respect to the C++standard, we enable this more efficient default transparent to the user. Preliminary experiments also confirm this in practice and further highlight that serialization incurs a non-negligible overhead, which is the reason why KaMPIng uses serialization only if explicitly enabled.
III-E Enabling Safety for Nonblocking Communication
Nonblocking communication in MPI is important for both correct and performant applications. MPI allows to initiate an operation, which returns a request handle to the user. A user than has to complete the request, by either testing for completion of the request using MPI_Test or using MPI_Wait to block until the request is completed. The semantics of MPI only allow updating send buffers or reading from receive buffers taking part in a previously initiated operation when the corresponding request has been completed. This introduces an additional source for programming errors, as MPI does not hinder users from accessing the memory locations regardless of completion status.
For asynchronous (I/O) operations the C++standard library provides std::future, which allows to query or wait for the result of an asynchronous operation, only returning a value once the operation has completed. Using std::future to provide a safe interface for nonblocking communicating is not possible, as they are tied to asynchronous progress happening in the background, which the MPI standard does not guarantee.
To solve this, we introduce a similar concept called a nonblocking MPI result, which encapsulates an MPI_Request and the data returned by value from the operation, as described in Section III-B. The data is only returned to the user by calling result.wait() which internally completes the request. Calling result.test() returns an std::optional, which only contains the returned data if the request is completed, and otherwise returns std::nullopt. This ensures that a user can only access valid received data.
To also prevent unwanted access to send buffers during nonblocking operations, the user can move the data into the call. The nonblocking result then assures that the data lives long enough and it is re-returned to the user upon completion of the call. This happens without any copying of data by relying on C++’s move semantics. See Figure 4 for an example.
@@ III-E Enabling Safety for Nonblocking Communication
Figure 4: Example of nonblocking safety in KaMPIng.
This is enabled by KaMPIng’s distinct parameter handling, making it the first C++MPI library which provides safety guarantees for nonblocking communication. Opposed to that, other MPI bindings offer no enhanced safety features for the data involved in a nonblocking call, but only return request handles. Here the user is still responsible for ensuring that no invalid data access happen while an operation is in progress. Only rsmpi provides similar guarantees, powered by Rust’s ownership model.
Another feature KaMPIng provides to facilitate working with nonblocking MPI are request pools, which allow for easy completion of multiple requests. The user only needs to submit the request associated with the call to such a pool. While the current implementation just collects them in an unbounded array, requests pool are designed with extensibility in mind, to enable more sophisticated variants. For example, we are currently working on a request pool with a fixed number of slots, internally maintaining free slots, which allows to limit the number of concurrent nonblocking requests.
@@ III-F Extensibility
While the main goal of KaMPIng is to design C++bindings for MPI which cover most usage scenarios, full coverage is impossible to achieve. We therefore designed it with extensibility and compatibility with existing MPI code in mind, to allow easy extension and alteration of its core features. KaMPIng’s plugin interface allows overriding member functions of a communicator object (e.g., collectives) and adding additional functionality without changing existing application code. The library allows plugin implementers to define new named parameters to enable the full named parameter flexibility also for these library extensions. This architecture allows us to keep KaMPIng’s core library small, while providing a base for third party general purpose MPI libraries and keeping maintenance low, in order not to follow the fate of the official MPI C++interface.
@@ III-G More Safety Features for MPI
KaMPIngtries to improve upon this by using three major techniques, following the C++core guidelines: using exceptions for failures [16, E.2], catching usage errors at compile time whenever possible [16, P.5], and making heavy use of assertions at runtime. While C++’s template-metaprogramming is notorious for complex and hard-to-read compiler errors, we try to ensure that compile time assertions fail early and provide helpful human-readable error messages. For example, when the user does not provide a required parameter to a collective operation, the error message indicates which parameter is missing during compile time. KaMPIng also includes many runtime assertions verifying MPI invariants, that are grouped in different levels, ranging from lightweight checks to assertions involving additional communication. The assertions can be completely disabled level-by-level. The use of exceptions can also be completley disabled, and KaMPIng allows to override the error handling strategy using the plugin system.
@@ IV Integrating KaMPIng into Real-World Applications
To highlight the usability of our library, we have integrated KaMPIng into multiple (research) applications, ranging from sorting (sample sort and suffix sorting) over graph algorithms (BFS and label propagation) to a large phylogenetic interference tool. Experiments backing our (near) zero-overhead are executed on up to 256 compute nodes of SuperMUC-NG, where each node is equipped with an Intel Skylake Xeon Platinum 8174 processor with $48$ cores. The internal interconnect is a fast OmniPath network with $100\text{\,}\mathrm{Gbit}\text{/}\mathrm{s}$. Our code is compiled with g++-12.2.0 and Intel MPI 2021 using optimization level -O3.
TABLE I: Lines of code for examples using KaMPIng vs. other bindings [^6] The source codes for all three examples is available at https://github.com/kamping-site/kamping-examples/..
@@ IV-A Sorting
As a first example, we use a textbook distributed sample sort [20] with Figure 5 showing a prototypical implementation in KaMPIng. We have implemented the sample sort algorithm using all previously discussed C++MPI bindings in a comparable way, where all shared parts of the code have been extracted to functions and the code has been formatted identically with clang-format using the Google style template. In this setting, we require only $16$ lines of code (LOC) using KaMPIng while the plain MPI example requires $32$ LOC. The lines of code for all bindings can be found in Footnote 6. Figure 6 shows the running time of the different implementations in a weak-scaling experiment. We sort a distributed array with $10^{6}$ 64-bit integers per rank, which are drawn uniformly at random.
We see that KaMPIng introduces no additional overhead compared to a hand-rolled implementation in plain MPI or other libraries, but makes the implementation a lot easier to read and to write, while being more flexible.
@@
For a more complex example we consider an application from text processing: We sort all suffixes of a text lexicographically, i.e., we compute the suffix array [9]. Here, we implemented two algorithms: DCX [21] and Prefix Doubling [9]. Our KaMPIng implementation of DCX requires 1 264 LOC whereas the plain MPI implementation [22] needs 1 396 LOC. The additional 9.5 % of code are mostly due to boilerplate code, e.g., distributing send counts for MPI_Alltoallv and the tedious construction of MPI types. Likewise, our KaMPIng implementation of the Prefix Doubling algorithm requires 163 LOC. An existing plain MPI implementation of the same algorithm [23] needs 426 LOC (not counting the 1 442 LOC for wrapped MPI functionality used by the plain MPI implementation). Even when using a high-level distributed programming framework, implementing the Prefix Doubling algorithm still requires 266 LOC [24].

### 23 Aug 2024 v2

@@ III-B Input and Output Parameters
Since one is primarily interested in the receive buffer in MPI calls, this parameter is always implicitly returned by KaMPIng. To retrieve other parameters from the wrapped MPI call one has to explicitly pass the corresponding out-parameter. This is a major improvement over previous MPI libraries, which simply mimic MPI’s C-Interface and return output data next to the receive buffer by pointer or reference. This is not in line with the C++ core guidelines which strongly suggest returning output data by value [9, F.20]. Furthermore, it is often unclear which additional parameters are computed by the library, as the overloaded wrapped MPI function calls differ only in the number of function arguments. Combined with the large number of parameters of MPI calls, it is hard to see which argument refers to which parameter when looking at the code.
@@
In the following, this is illustrated with a call to the wrapped MPI_Allgatherv function.
@@
A shortcoming of returning by value is the redundant memory allocation in case a previously allocated container could be reused instead. In KaMPIng, we offer two solutions for this scenario. For containers supporting C++ move semantics, a previously user-allocated container can be simply moved to the underlying call and is then subsequently returned with the result object by value.
@@
If there is no efficient way to support move semantics for a container type, it is also possible to pass the container via reference to the underlying call. The data computed by KaMPIng for the requested parameter will then be written directly to the specified memory location.
@@ III-C Controlling Memory Allocation
Previous MPI wrappers have no unified way of controlling memory allocation. They either accept containers that are always resized to fit, or, if resizing is not desired, the user has to pass raw pointers directly. They also offer no control over allocation happening for default parameter computation.
@@
If KaMPIng has to create auxiliary data structures to compute missing parameters, the user may either pass preallocated containers to use or provide the container’s type via template parameters. Recall that additional allocation is omitted entirely when parameters are provided by the user.
As all of this relies on template metaprogramming, there is no additional overhead compared to a hand-rolled implementation. This flexibility allows to quickly implement distributed algorithms and then iteratively fine-tune memory allocations and library inferred values. This facilitates an algorithm engineering workflow that involves iterative refinement of implementations and analysis through experimentation.
@@ III-D Using Custom Types
HPC applications use a variety of data types that need to be communicated. Beyond basic data types corresponding directly to C++’s built-in types, MPI allows for complex derived data types using type constructors, such as MPI_Type_create_struct and MPI_Type_create_contiguous.
C’s lack of type introspection forces users to always pass the type explicitly to a communication call, which is both tedious and error-prone, as type definitions need to be kept in sync with the actual data layout.
@@ III-D1 Static derived data types at compile time
namespace kamping {
// using KaMPIng’s built-in struct serializer
@@
Existing MPI bindings also support non-built-in static types to some extent. For Boost.MPI that is dependent on the definition of a serialization function. Similar to KaMPIng, types are managed using a global type pool, but each operation incurs a runtime type lookup. MPL and RWTH-MPI use the Construct-On-First-Use-Idiom to commit types before first use, and MPL also uses PFR to provide automatic type definitions by using reflection. Opposed to KaMPIng, types are not properly freed, which may result in resource leakage. RWTH-MPI allows custom static type definitions, but the user is responsible for committing and freeing types.
@@ III-D2 Support for dynamic types
In KaMPIng, dynamic types can currently only be constructed using MPI’s type constructors. MPL on the other hand provides a runtime type interface mirroring MPI’s type constructors using the builder pattern, which encapsulates the constructed types in so-called layouts. While MPL’s collective operations are tightly interleaved with the layout system, which results in verbose code, its type construction is a powerful feature that we plan to integrate as the default way for constructing dynamic types in KaMPIng. RWTH-MPI offers no support for dynamic types.
@@ III-D3 Communicating arbitrary data using serialization
Some applications require communicating non-contiguous data which is (partially) located on the heap, e.g., std::string or std::unordered_map and cannot be represented using MPI datatypes. These have to be packed into a contiguous buffer before communication. KaMPIng facilitates this by providing serialization support, which is both highly tuneable and transparent to the user. We rely on the popular C++ serialization library Cereal [21], which supports STL data types and allows providing serialization routines for custom types. While serialization is transparent to the user, i.e., the user never sees the serialized data, it still has to be explicitly enabled as it usually incurs hidden costs for allocating memory for serialized data and performing (de-)serialization. Through Cereal’s flexible design, serialization in KaMPIng is also highly configurable; allowing users to specify custom serialization functions and archives, e.g., binary formats, JSON, or XML. See Figure 5 for an example of how to use serialization in KaMPIng.
@@
recv_buf(as_deserializable<dict>())
@@
Besides KaMPIng, Boost.MPI is the only library offering serialization support, but is tightly coupled with other Boost libraries and serialization is performed implicitly; if a type is not marked as an MPI data type, serialization is used. This makes it hard to see whether costly serialization is involved by just looking at the code.
@@ III-D4 Towards sensible defaults for type construction
By default, KaMPIng maps trivially copyable types to a type interpreted as a contiguous sequence of bytes. When defining a struct type where the members have alignment gaps, MPI does not include the gap in the communicated data. This requires non-contiguous memory accesses, which may be slower than copying whole memory blocks to the communication hardware (MPI standard [1, §5.1.6]). The standard suggests introducing dummy struct members to fill these gaps, but this requires the user to modify their non-MPI code. By using a type consisting of contiguous bytes when valid with respect to the C++ standard, we enable this more efficient default transparent to the user. Preliminary experiments also confirm this in practice and further highlight that serialization incurs a non-negligible overhead, which is the reason why KaMPIng uses serialization only if explicitly enabled.
III-E Enabling Safety for Non-blocking Communication
Non-blocking communication in MPI is important for both correct and performant applications. MPI allows to initiate an operation, which returns a request handle to the user. A user then has to complete the request, by either testing for completion of the request using MPI_Test or using MPI_Wait to block until the request is completed. The semantics of MPI only allow updating send buffers or reading from receive buffers taking part in a previously initiated operation when the corresponding request has been completed. This introduces an additional source for programming errors, as MPI does not hinder users from accessing the memory locations regardless of completion status.
For asynchronous (I/O) operations the C++ standard library provides std::future, which allows querying or waiting for the result of an asynchronous operation, only returning a value once the operation has completed. Using std::future to provide a safe interface for non-blocking communicating is not possible, as they are tied to asynchronous progress happening in the background, which the MPI standard does not guarantee.
To solve this, we introduce a similar concept called a non-blocking MPI result, which encapsulates an MPI_Request and the data returned by value from the operation, as described in Section III-B. The data is only returned to the user by calling result.wait() which internally completes the request. Calling result.test() returns an std::optional, which only contains the returned data if the request is completed, and otherwise returns std::nullopt. This ensures that a user can only access valid received data.
To also prevent unwanted access to send buffers during non-blocking operations, the user can move the data into the call. The non-blocking result then assures that the data lives long enough, and it is re-returned to the user upon completion of the call. This happens without any copying of data by relying on C++’s move semantics. See Figure 6 for an example.
@@ III-E Enabling Safety for Non-blocking Communication
Figure 6: Example of non-blocking safety in KaMPIng.
This is enabled by KaMPIng’s distinct parameter handling, making it the first C++ MPI library that provides safety guarantees for non-blocking communication. Opposed to that, other MPI bindings offer no enhanced safety features for the data involved in a non-blocking call, but only return request handles. Here the user is still responsible for ensuring that no invalid data access happens while an operation is in progress. Only rsmpi provides similar guarantees, powered by Rust’s ownership model.
Another feature KaMPIng provides to facilitate working with non-blocking MPI are request pools, which allow for easy completion of multiple requests. The user only needs to submit the request associated with the call to such a pool. While the current implementation just collects them in an unbounded array, requests pool are designed with extensibility in mind, to enable more sophisticated variants. For example, we are currently working on a request pool with a fixed number of slots, internally maintaining free slots, which allows limiting the number of concurrent non-blocking requests.
@@ III-F Extensibility
While the main goal of KaMPIng is to design C++ bindings for MPI which can be used in any MPI application, designing a one-size-fits-all library is hard. To enable further flexibility we therefore designed it with extensibility and compatibility with existing MPI code in mind, to allow easy extension and alteration of its core features. On the one hand, an existing code base can be gradually migrated to use only some parts of KaMPIng, because it is fully compatible with native MPI objects, such as request or type handles. On the other hand, KaMPIng’s plugin interface allows overriding member functions of a communicator object (e.g., collectives) and adding additional functionality without changing existing application code. The library allows plugin implementers to define new named parameters to enable the full named parameter flexibility also for these library extensions. This architecture allows us to keep KaMPIng’s core library small while providing a base for third-party general-purpose MPI libraries and keeping maintenance low, in order not to follow the fate of the official MPI C++ interface.
@@ III-G More Safety Features for MPI
In addition to the ease of use and safety features introduced by KaMPIng’s parameter handling, it also provides other means of preventing sources of errors in MPI code.
Error handling
@@
KaMPIngtries to improve upon this by using three major techniques, following the C++ core guidelines: using exceptions for failures [9, E.2], catching usage errors at compile time whenever possible [9, P.5], and making heavy use of assertions at runtime. While C++’s template metaprogramming is notorious for complex and hard-to-read compiler errors, we try to ensure that compile-time assertions fail early and provide helpful human-readable error messages, e.g., for missing parameters or incompatible types. For example, when the user does not provide a required parameter to a collective operation, the error message indicates which parameter is missing during compile time. KaMPIng also includes many runtime assertions verifying MPI invariants, that are grouped in different levels, ranging from lightweight checks to assertions involving additional communication. The assertions can be completely disabled level-by-level. The use of exceptions can also be completely disabled, and KaMPIng allows overriding the error handling strategy using the plugin system.
@@
Simplified MPI_IN_PLACE
KaMPIngalso tries to prevent programming errors when working with in-place MPI operations. Consider using the in-place variant of MPI_Allgather, as in Figure 2: Using MPI’s C interface, the user has to explicitly pass MPI_IN_PLACE on all ranks as send buffer and the send count and type parameters are ignored. KaMPIng’s concept of in-out-parameters simplifies the call semantics of in-place calls. If a user passes data as send_recv_buf instead of send_buf, then KaMPIng automatically passes the correct arguments to the underlying in-place call and issues a compilation error if the user provides an argument which would be ignored by the in-place call. This is also compatible with move semantics, allowing for concise in-place calls as shown below:
std::vector<int> data(comm.size());
data[comm.rank()] = ...;
data = comm.allgatherv(send_recv_buf(std::move(data)));
III-H Implementation Details
KaMPIngmakes extensive use of C++’s template metaprogramming capabilities. For example, using constexpr if and SFINAE [22, 8.4] (Substitution Failure is Not an Error), often allows us to eliminate runtime control flow constructs in the wrappers around MPI calls, preventing costly branch mispredictions. Named parameters are implemented using template parameter packs of parameter objects. Compile-time checks for the absence of each parameter then allow to conditionally enable logic for computing default values via constexpr if.
All containers passed as arguments are wrapped in a templated DataBuffer which handles ownership and modifiability transparently at compile time [^4] Because this works with any STL-compliant container, KaMPIng also supports accelerator-aware MPI implementations directly. Pointers or containers (like thrust::device_vector) to device memory can be passed just like normal containers.. These data buffers are moved to a templated result object, which is returned to the user and can be destructed using structured bindings as described in Section III-B. As we do not copy but move the data, this imposes nearly zero overhead.
All wrapped MPI functionality has been extensively tested using a large number of parameter combinations as part of our development cycle. All (unit-)tests are executed using the C++ compilers clang and gcc (in debug and release mode) and OpenMPI. We use MPI’s profiling interface to ensure that only the expected MPI calls are issued if KaMPIng calls MPI internally to compute default values. Additionally, we evaluated our benchmarks using OpenMPI and IntelMPI on the supercomputers SuperMUC-NG and HoreKa, as detailed in the following section.
@@ IV Integrating KaMPIng into Real-World Applications
To highlight the usability of our library, we have integrated KaMPIng into multiple (research) applications, ranging from sorting (sample sort and suffix sorting) over graph algorithms (BFS and label propagation) to a large phylogenetic interference tool. Experiments backing our (near) zero overhead claim are executed on up to 256 compute nodes of SuperMUC-NG [^5] We also conducted some experiments on the smaller HoreKa supercomputer. As the results obtained there are similar to our findings from SuperMUC-NG, we omit them here., where each node is equipped with an Intel Skylake Xeon Platinum 8174 processor with $48$ cores. The internal interconnect is a fast OmniPath network with $100\text{\,}\mathrm{Gbit}\text{/}\mathrm{s}$. Our code is compiled with g++-12.2.0 and Intel MPI 2021 using optimization level -O3.
TABLE I: Lines of code for examples using KaMPIng vs. other bindings [^7] The source codes for all three examples are available at https://github.com/kamping-site/kamping-examples/..
@@ IV-A Sorting
As a first example, we use a textbook distributed sample sort [23] with Figure 7 showing a prototypical implementation in KaMPIng. We have implemented the sample sort algorithm using all previously discussed C++ MPI bindings comparably, where all shared parts of the code have been extracted to functions and the code has been formatted identically with clang-format using the Google style template. In this setting, we require only $16$ lines of code (LOC) using KaMPIng while the plain MPI example requires $32$ LOC. The lines of code for all bindings can be found in Footnote 7. Figure 8 shows the running time of the different implementations in a weak-scaling experiment. We sort a distributed array with $10^{6}$ 64-bit integers per rank, which are drawn uniformly at random.
We see that KaMPIng introduces no additional overhead compared to a hand-rolled implementation in plain MPI or other libraries, but makes the implementation a lot easier to read and write while being more flexible.
@@
For a more complex example, we consider an application from text processing: We sort all suffixes of a text lexicographically, i.e., we compute the suffix array [12]. Here, we implemented two algorithms: DCX [24] and Prefix Doubling [12]. Our KaMPIng implementation of DCX requires 1 264 LOC whereas the plain MPI implementation [25] needs 1 396 LOC. The additional 9.5 % of code is mostly due to boilerplate code, e.g., distributing send counts for MPI_Alltoallv and the tedious construction of MPI types. Likewise, our KaMPIng implementation of the Prefix Doubling algorithm requires 163 LOC. An existing plain MPI implementation of the same algorithm [26] needs 426 LOC (not counting the 1 442 LOC for wrapped MPI functionality used by the plain MPI implementation). Even when using a high-level distributed programming framework, implementing the Prefix Doubling algorithm still requires 266 LOC [27].